Find the angles of an isosceles triangle whose third angle is ten degrees less than three times the base angles.
Answer by chessace(471)

A3 = 3 * A1 - 10
A3 + 2 * A1 = 180
-2A1: A3 = -2 * A1 + 180
Subtract line 3 from line 1: 0 = (3 - (-2)) * A1 - 10 - 180
5 * A1 - 190 = 0
A1 = 38
A3 = 3*38-10 = 114-10 = 104
Double check: 2(38) + 104 = 76+104 = 180
A1=38, A2=38, A3=104